What are the dimensions of a rectangle that has a perimeter of 74 units and an area of 102 square units?

Answer by ewatrrr(24785)

 
Hi,
.
L + w = 37
L = 37-w
A = Lw = (37-w)w = 102
w^2 -37w + 102 = 0
(w-34)(w-3) = 0
Dimensions are 34units by 3units